Required information [The following information applies to the questions displayed below.] Tyrell Co. entered into the following transactions involving short-term liabilities in 2017 and 2018. 2017 Apr. 20 Purchased $35,500 of merchandise on credit from Locust, terms n/30. Tyrell uses the perpetual inventory system. May 19 Replaced the April 20 account payable to Locust with a 90-day, $35,000 note bearing 8% annual interest along with paying $500 in cash. July 8 Borrowed $63,000 cash from NBR Bank by signing a 120-day, 11% interest-bearing note with a face value of $63,000. __?__ Paid the amount due on the note to Locust at the maturity date. __?__ Paid the amount due on the note to NBR Bank at the maturity date. Nov. 28 Borrowed $33,000 cash from Fargo Bank by signing a 60-day, 9% interest-bearing note with a face value of $33,000. Dec. 31 Recorded an adjusting entry for accrued interest on the note to Fargo Bank. 2018 __?__ Paid the amount due on the note to Fargo Bank at the maturity date.

5.1 Prepare journal entries for all the preceding transactions and events for 2017. (Do not round your intermediate calculations.)
Answer #1
1) Maturity date
locust NBR fargo
date of the note 19-May 8-Jul 28-Nov
term of note 90 120 60
maturity date 17-Aug 5-Nov 27-Jan
2) interest due at maturity
principal * Rate * time = interest
locust 35,000 * 8% * 90/360 = 700
NBR 63,000 * 11% * 120/360 = 2310
Fargo 33,000 * 9% * 60/360 = 495
3) Amount in adjusting entry
Fargo Bank
principal * Rate * time = interest
interest to be acccrued in 2016 33,000 * 9% * 33/360 = 272
4) interest expense to be recorded in 2017
principal * Rate * time = interest
interest to recorded in 2018 33,000 * 9% * 27/360 = 223
Journal entries
Date Accounting titles & Explanations Debit Credit
2016
20-Apr inventory 35,500
Accounts payable 35,500
19-May Accounts payable 35,500
cash 500
notes payable 35,000
8-Jul Cash 63,000
notes payable 63,000
17-Aug notes payable 35,000
interest expense 700
cash 35,700
5-Nov notes payable 63,000
interest expense 2,310
cash 65,310
28-Nov Cash 33,000
notes payable 33,000
31-Dec interest expense 272
interest payable 272
2017
27-Jan notes payable 33,000
interest payable 272
interest expense 223
cash 33,495